OpenStudy (anonymous):

What is the first step that you have to do in order to solve the proportion: x/2=4/12 A. 8.5 B. 26y = 3 C. y = 3(26) D. 3(y) = 1(26)

OpenStudy (anonymous):

so what would the answer be

OpenStudy (ivettef365):

why you have an x on the proportion and a y on the answers??

OpenStudy (anonymous):

becuz these are answer choices

OpenStudy (ivettef365):

those answer choices makes no sense, are you sure you have the right problem for those answer choices

OpenStudy (anonymous):

yea im sure

OpenStudy (ivettef365):

let see if somebody else makes sense of it, because proportion is cross multiplying you have the problem as x 4 ---- = ---- 2 12 that is equal to 12(x) = 4(2) 12x = 8 x = 8/12 x = 2/3